### v3

@@ Abstract
The recent development of quantum computing, which uses entanglement, superposition, and other quantum fundamental concepts, can provide substantial processing advantages over traditional computing. These quantum features help solve many complex problems that cannot be solved with conventional computing methods. These problems include modeling quantum mechanics, logistics, chemical-based advances, drug design, statistical science, sustainable energy, banking, reliable communication, and quantum chemical engineering. The last few years have witnessed remarkable advancements in quantum software and algorithm creation and quantum hardware research, which has significantly advanced the prospect of realizing quantum computers. It would be helpful to have comprehensive literature research on this area to grasp the current status and find outstanding problems that require considerable attention from the research community working in the quantum computing industry. To better understand quantum computing, this paper examines the foundations and vision based on current research in this area. We discuss cutting-edge developments in quantum computer hardware advancement and subsequent advances in quantum cryptography, quantum software, and high-scalability quantum computers. Many potential challenges and exciting new trends for quantum technology research and development are highlighted in this paper for a broader debate.
@@
Quantum Computing , Artificial Intelligence , Machine Learning , Cryptography , Cyber Security , Qubits
@@ 1 Promising Age of Quantum Computing
However, quantum computing is one of the most talked-about fields right now (as of 2024), and it has been growing at a tremendous pace in recent years [6]. There is a great deal of enthusiasm among academics and businesses alike to construct initial quantum computers due to their promise of providing, for certain tasks, processing powers beyond those of current most powerful supercomputers. Strong efforts to build large-scale quantum computers are now underway with several established corporations (Chinese companies like ZTE, QUDOOR and USA based companies such as Honeywell, Intel, Google, Microsoft, and IBM), growing small and medium-sized enterprises (e.g., D-Wave), and aspiring startups (e.g., Rigetti, Xanadu, Infleqtion, Origin Quantum, and IonQ). There has been enormous advancement in quantum algorithms and quantum software in recent years, which has occurred in tandem with the development of quantum hardware.
@@
It is well-known that traditional digital computing relies on bits that are limited to two possible values—‘0’ or ‘1’—to store and process data. In quantum computing, the corresponding unit is the quantum bit (qubit) that, according to quantum physics, may have either a value of ‘0’ or ‘1’ or exist on a superposition of the two (functionally being in both states simultaneously!) [7, 8, 9]. Because of this, quantum computers have access to a computational field (known as Hilbert space [10]) of huge dimension, where $n$ qubits might be in a superposition state with $2^{n}$ potential values at any one moment. Due to the exponential growth of the parameter space, problems on a large scale are expected to be easier to solve with the advent of quantum computers. Nevertheless, developing a large-scale quantum computer has its own set of specific challenges. The most demanding to mitigate is the decoherence of the quantum states on which qubits are encoded. Decoherence happens when qubits interact with their surroundings and lose their coherent features. For that it represents one of the biggest obstacles to developing large-scale quantum devices [11]. Assuming the unavoidable presence of environmental noise, “Noisy Intermediate Scale Quantum (NISQ)” devices, try to deal with imperfections and losses driven by decoherence. Reducing the probability of decoherence and creating effective error correction procedures to overcome defects in NISQ devices are important goals of current studies in quantum computing [12]. The second big problem with modern quantum devices is to identify approaches to effectively engineer and interconnect qubits [13]. At the moment of writing current quantum devices are able to deal with relatively sparsely connected qubits, making it difficult to map deep quantum circuits with multiple two-qubit gates that necessitate strong couplings between qubits [14].
@@
1.1 Quantum Supremacy: Regardless of technological hurdles, NISQ quantum computers have shown promising computing capability in their early stages. Google’s recent proof of quantum supremacy is a major step forward for quantum computing [15]. There is currently a worldwide race to be the first to implement quantum computing in order to tackle a practical problem that a conventional computer cannot solve in a reasonable time — also known as “quantum advantage”. To reach this desired level of quantum computing, it is necessary to reduce the probability of the decoherence of qubits drastically through improvements in quantum hardware, quantum algorithms, and error correction during the upcoming years. A lot of work is being put into developing and benchmarking quantum algorithms using NISQ devices. While Grover’s and Shor’s quantum algorithms were among the first that stood out in the early 1990s, hundreds of other algorithms have been invented since then. Variational Quantum Eigensolver (VQE) [16, 17] and other variational quantum algorithms [18] are a popular kind of hybrid quantum-classical algorithm that combines the advantages of the two technologies. On NISQ devices, VQE algorithms have performed very well in solving quantum mechanical problems and Quantum Artificial Intelligence (QAI) tasks [19]. While a large and resilient quantum computer is not available yet and will still require significant improvements before its full promise for practical applications can be realised, quantum computing is already available for research and prototype scenarios with encouraging results on current NISQ-era equipment [20].
@@
When applied to classical data, QAI has the potential to greatly accelerate machine intelligence techniques [21, 22]. Quantum neural networks, quantum support vector machines, and quantum principle component evaluation have been studied [23, 24], and some recent research returned encouraging findings [25], although it is still not completely known if quantum neural networks will provide better computing efficiency than traditional machine learning executions.
@@
Measurement is the last stage of a quantum computation; it collapses the stochastic quantum state into a deterministic state. Although quantum algorithms typically guarantee that the correct outcome has the highest likelihood, the stochastic nature of the process cannot guarantee that the correct outcome is actually sampled. Therefore, some classical post-processing (such as majority voting or statistical estimation) is usually needed to produce a final output from the raw results obtained with the quantum computer.
@@ 3 Technological Advances and Software Tools
The invention of quantum software is an emerging yet relatively less developed subject compared to quantum modelling and quantum technology [36]. Several quantum applications are already accessible from various platforms/sources, including Google, IBM, Microsoft, and D-Wave. Quantum programming tools have been produced at an increasing pace; however, there is a lack of excellent programming tools, similar to conventional programming languages like C++ and Java, and these applications are still rather low-level, like assembly-level languages. A number of important areas have been identified in recent research pertaining to software programs that use quantum computing, including coding languages, programmers, error-correction firmware, physical level schedulers and optimisers, logical level schedulers and optimisation techniques, and hardware control of software updates. Most important topics to study in the field are [37, 38, 39]: (i) frameworks, semantics and compilation of programming language; (ii) workflows, controlled and adjoint operations & clean and borrowed qubits and (iii) simulators. Integrating quantum algorithms with defective equipment effectively is the goal of powerful quantum error-correcting firmware [37]. Located at the very bottom of the quantum computing stack, error-correcting quantum firmware aids in lowering the error rate due to flawed hardware, as well as the intricacy and resource consumption of the system [38]. It is envisaged that software managing quantum hardware would have outstanding performance, be able to use sophisticated quantum management techniques, have top-quality effects at the system level, be able to regulate for both global and local optimal outcomes through simulation, and have adequate physical schedules [39]. At this date, notwithstanding the absence of a single programming framework/model able to overcome the others, there are different platforms for quantum computer programming, often provided and “tied” to the provided hardware solutions. Among the most famous are: Qiskit (Quantum Information Science Kit) — developed by IBM [^1] https://www.ibm.com/quantum/qiskit; Cirq — developed by Google [^2] https://quantumai.google/cirq; and PyQuil — developed by Rigetti Computing [^3] https://github.com/rigetti/pyquil. To push their solutions, quantum developers often release these frameworks with open-source licenses and with an Application Programming Interface (API) in Python, which is a language that is straightforward to learn. Quantum Annealing is following the same path, with a couple of “programming frameworks” — i.e., D-Wave Ocean Software and Leap — both provided by D-Wave. Recently, Fujitsu’s Digital Annealer has been promising to bring quantum-inspired technology using traditional computing platforms [40].
@@ 4 Modern Cryptography: From Quantum to Post-Quantum
The advent of quantum computers heralds the ground-breaking era within the realm of data integrity and cybersecurity. With scalable computing power, quantum computers can effortlessly break the security of traditional cryptosystems. Traditional cryptosystems rely on factorization and discrete logarithms, both of which are considered hard problems for classical computers. However, quantum computers have efficient processing capabilities to solve these hard problems within polynomial time [41]. For example, an adversary can easily break the RSA (Rivest-Shamir-Adleman) security in polynomial time by exploiting Shor’s algorithm for factoring large numbers on a quantum computer, posing potential threats to the integrity of communication networks  [42]. This potential breakthrough has led to the development of quantum cryptography based on the principles of quantum physics. Quantum cryptography uses polarization and bits designations methods to exchange data, where it first converts the data into bits and transfers the converted data into photons. Transmission of quantum keys, encryption/decryption, signatures, authentication, and hashing are all examples of cryptographic operations that fall under the umbrella term “quantum cryptography” [41, 43]. Instead of using conventional algorithms founded on arithmetic that uses fewer resources and is easier to use, quantum security makes use of the basic quantum principles (entanglement) [44]. Furthermore, since it is impossible to recreate the exact quantum state of a set of qubits and the information encoded within it, it is significantly more difficult for an adversary to develop an attack since they would require continuous access [45]. This also simplifies the implementation of snooping detection. Further, quantum-based principles make novel approaches to network design and architecture possible. This approach, known as quantum internet, promises security improvements through quantum authenticity, long- and short-distance entanglement, improved data analysis driven by quantum computing accelerated machine learning (ML), and quantum-specific anomaly detection. Some of the methods used for anomaly detection focus on recognizing efficiency mismatches, detector blinding, detection system dead-time, beam-splitter, spatial mode, and eavesdropping [46].
@@
4.1 Quantum Key Distribution: Since quantum algorithms (e.g., Shor’s) endanger the classical cryptographic key exchange techniques (e.g., Diffie–Hellman key exchange), the Quantum Key Distribution (QKD) has been developed as a potential remedy to resist quantum attacks. QKDs protocols are always hybrid, i.e., they rely on both quantum and classical communications to implement a virtually impenetrable crypto-system that promises to protect the privacy of communication even against attacks conducted by quantum computers. An essential tenet of such an approach to encryption is roothed in quantum physics and takes the shape of the no-cloning theorem [47], which asserts that a flawless replica of arbitrary (i.e. non-orthogonal) quantum states cannot be created without corrupting the probed quantum states. That means that any effort to gain insight into such non-orthogonal quantum states, used to encode information, would results in the introduction of noise that can be readily identified by either the key distributor or the recipient (the parties, conventionally Alice and Bob). Such a mechanism allows the parties to quantify the amount of information potentially eavesdropped during the quantum communication, and then apply classical protocols of error correction and privacy amplification that can reduce the eavesdropper’s knowledge on the shared key to a negligible amount. This second part of the cryptosystem is usually called classical communication phase phase.
@@
4.2 Post-Quantum Cryptography: Conventionally, the security of classical cryptographic primitives (e.g., RSA, Diffie–Hellman, etc.) depends on the hard problems of discrete arithmetic, prime factorization of integers, and elliptic-curve discrete logarithms. Sadly, these present-day cryptographic primitives based on such hard problems might theoretically be solvable in a brief span of time using the possible applications of quantum computers. The imminent attacks of quantum algorithms posed on conventional cryptographic protocols have promoted a sense of urgency in designing alternative schemes to mitigate quantum attacks. Such alternatives are generally characterized as post-quantum cryptography (PQC). These schemes can effectively deal with prevalent challenges triggered by quantum adversaries. This leads many to wonder about post-quantum cryptography [48]. The protocols used in post-quantum cryptography may generally be grouped into five types: code-based, hash-based, lattice-based, multifaceted, and supersingular curve-elliptic isogeny schemes [11].
Threatening the security of encrypted defense data, quantum computing’s greater processing capabilities expose prospective shortcomings in present encryption techniques. The “harvest now, decrypt later” approach is authentic since it allows attackers to store encrypted material until decryption technology is more advanced [49]. Since the seriousness of its failures determines how willing it is to take risks, the military industry cannot afford to ignore or minimize risk.

### v4

@@ Abstract
The recent development of quantum computing, which uses entanglement, superposition, and other quantum fundamental concepts, can provide substantial processing advantages over traditional computing. These quantum features help solve many complex problems that cannot be solved otherwise with conventional computing methods. These problems include modeling quantum mechanics, logistics, chemical-based advances, drug design, statistical science, sustainable energy, banking, reliable communication, and quantum chemical engineering. The last few years have witnessed remarkable progress in quantum software and algorithm creation and quantum hardware research, which has significantly advanced the prospect of realizing quantum computers. It would be helpful to have comprehensive literature research on this area to grasp the current status and find outstanding problems that require considerable attention from the research community working in the quantum computing industry. To better understand quantum computing, this paper examines the foundations and vision based on current research in this area. We discuss cutting-edge developments in quantum computer hardware advancement and subsequent advances in quantum cryptography, quantum software, and high-scalability quantum computers. Many potential challenges and exciting new trends for quantum technology research and development are highlighted in this paper for a broader debate.
@@
[…]
@@ 1 Promising Age of Quantum Computing
However, quantum computing is one of the most talked-about fields (as of 2024), and its progress has been growing at a tremendous pace in recent years [6]. There is a great deal of enthusiasm among academics and businesses alike to construct initial quantum computers due to their promise of providing, for certain tasks, processing powers beyond those of current most powerful supercomputers. Strong efforts to build large-scale quantum computers are now underway with several established corporations (Chinese companies like ZTE, QUDOOR and USA based companies such as Honeywell, Intel, Google, Microsoft, and IBM), growing small and medium-sized enterprises (e.g., D-Wave), and aspiring startups (e.g., Rigetti, Xanadu, Infleqtion, Origin Quantum, and IonQ). There has been enormous advancement in quantum algorithms and quantum software in recent years, which has occurred in tandem with the development of quantum hardware.
@@
It is well-known that traditional digital computing relies on bits that are limited to two possible values—‘0’ or ‘1’—to store and process data. In quantum computing, the corresponding unit is the quantum bit (qubit) that, according to quantum physics, may have either a value of ‘0’ or ‘1’ or exist on a superposition of the two (functionally being in both states simultaneously!) [7, 8, 9]. Because of this, quantum computers have access to a computational field (known as Hilbert space [10]) of huge dimension, where $n$ qubits might be in a superposition state with $2^{n}$ potential values at any one moment. Due to the exponential growth of the parameter space, problems on a large scale are expected to be easier to solve with the advent of quantum computers. Nevertheless, developing a large-scale quantum computer has its own set of challenges. The most demanding to mitigate is the decoherence of the quantum states on which qubits are encoded. Decoherence happens when qubits interact with their surrounding environment and lose their coherent features. For that it represents one of the biggest obstacles to developing large-scale quantum devices [11]. Assuming the unavoidable presence of environmental noise, “Noisy Intermediate Scale Quantum (NISQ)” devices, try to deal with imperfections and losses driven by decoherence. Reducing the probability of decoherence and creating effective error correction procedures to overcome defects in NISQ devices are important goals of current studies in quantum computing [12]. The second big problem with modern quantum devices is to identify approaches to effectively engineer and interconnect (entangle) qubits [13]. At the moment of writing current quantum devices are able to deal with relatively sparsely connected qubits, making it difficult to map deep quantum circuits with multiple two-qubit gates that necessitate strong couplings between qubits [14].
@@
1.1 Quantum Supremacy: Regardless of technological hurdles, NISQ quantum computers have shown promising computing capability in their early stages. Google’s recent proof of quantum supremacy is a major step forward for quantum computing [15]. There is currently a worldwide race to be the first to implement quantum computing in order to tackle a practical problem that a conventional computer cannot solve in a reasonable time — also known as “quantum advantage”. To reach this desired level of quantum computing, it is necessary to reduce the probability of the decoherence of qubits drastically through improvements in quantum hardware, quantum algorithms, and error correction during the upcoming years. A lot of work is being put into developing and benchmarking quantum algorithms using NISQ devices. While Shor’s and Grover’s quantum algorithms were among the first that stood out in the early 1990s, hundreds of other algorithms have been invented since then. Variational Quantum Eigensolver (VQE) [16, 17] and other variational quantum algorithms [18] are a popular kind of hybrid quantum-classical algorithm that combines the advantages of the two technologies. On NISQ devices, VQE algorithms have performed exceptionally well in solving quantum mechanical problems and Quantum Artificial Intelligence (QAI) tasks [19]. While a large and resilient quantum computer is not available yet and will still require significant advancements before its full promise for practical applications can be realised, quantum computing is already available for research and prototyping scenarios with encouraging results on current NISQ-era equipment [20].
@@
When applied to classical data, QAI has the potential to greatly accelerate machine intelligence techniques [21, 22]. Quantum neural networks, quantum support vector machines, and quantum principle component evaluation have been studied [23, 24], and some recent research returned encouraging findings [25], although it is still not completely known if quantum neural networks will provide better computing efficiency than traditional machine learning techniques.
@@
Measurement is the last stage of a quantum computation; it collapses the stochastic quantum state into a deterministic state. Although quantum algorithms typically guarantee that the correct outcome has the highest likelihood, the stochastic nature of the process cannot guarantee that the correct outcome is actually sampled. Therefore, some classical post-processing (such as majority voting or statistical estimation) or repeating the computation several times is usually needed to produce a final output from the raw results obtained with the quantum computer.
@@ 3 Technological Advances and Software Tools
The invention of quantum software is an emerging yet relatively less developed field compared to quantum modelling and quantum technology [36]. Several quantum applications are already accessible from various platforms/sources, including Google, IBM, Microsoft, and D-Wave. Quantum programming tools have been produced at an increasing pace; however, there is a lack of excellent programming tools, similar to conventional programming languages like C++ and Java, and these applications are still rather low-level, like assembly-level languages. A number of important areas have been identified in recent research pertaining to software programs that use quantum computing, including coding languages, programmers, error-correction firmware, physical level schedulers and optimisers, logical level schedulers and optimisation techniques, and hardware control of software updates. Most important topics to study in the field are [37, 38, 39]: (i) frameworks, semantics and compilation of programming language; (ii) workflows, controlled and adjoint operations & clean and borrowed qubits and (iii) simulators. Integrating quantum algorithms with defective equipment effectively is the goal of powerful quantum error-correcting firmware [37]. Located at the very bottom of the quantum computing stack, error-correcting quantum firmware aids in lowering the error rate due to flawed hardware, as well as the intricacy and resource consumption of the system [38]. It is envisaged that software managing quantum hardware would have outstanding performance, be able to use sophisticated quantum management techniques, have top-quality effects at the system level, be able to regulate for both global and local optimal outcomes through simulation, and have adequate physical schedules [39]. At this date, notwithstanding the absence of a single programming framework/model able to overcome the others, there are different platforms for quantum computer programming, often provided and “tied” to the provided hardware solutions. Among the most famous are: Qiskit (Quantum Information Science Kit) — developed by IBM [^1] https://www.ibm.com/quantum/qiskit; Cirq — developed by Google [^2] https://quantumai.google/cirq; and PyQuil — developed by Rigetti Computing [^3] https://github.com/rigetti/pyquil. To push their solutions, quantum developers often release these frameworks with open-source licenses and with an Application Programming Interface (API) in Python, which is a language that is straightforward to learn. Quantum Annealing is following the same path, with a couple of “programming frameworks” — i.e., D-Wave Ocean Software and Leap — both provided by D-Wave. Recently, Fujitsu’s Digital Annealer has been promising to bring quantum-inspired technology using traditional computing platforms [40].
@@ 4 Modern Cryptography: From Quantum to Post-Quantum
The advent of quantum computers heralds a new ground-breaking era within the realm of data integrity and cybersecurity. With improving scalable computing power, quantum computers can effortlessly break the security of traditional cryptosystems, relying on factorization and discrete logarithms, both of which are considered hard problems for classical computers. By constrast, quantum computers have efficient processing capabilities to solve these hard problems within polynomial time [41]. For example, an adversary equipped with a quantum computer may break the RSA (Rivest-Shamir-Adleman) security in polynomial time by exploiting Shor’s algorithm for factoring large numbers. It is clear that such a possibility, despite not yet practical, poses potential threats to the integrity of communication networks  [42] that need to be analyzed and mitigated. In fact the potential threat represented by the Shor’s algorithm has led to new developments in classical cryptographic approaches, with the work on post-quantum cryptography and on a completely new paradigm to grant security named quantum cryptography  [43], or more precisely Quantum-Key Distribution (QKD). The novelty of QKD is that, instead of adding layers of security based on conventional (i.e. computationally hard to solve) algorithms, it uses fundamental properties of quantum particles to protect information from unauthorized parties. QKD protocols, are themselves composite algorithms where transmission of quantum signals, encryption/decryption, signatures, authentication, and hashing are all combined  [44] to achieve (theoretically) unconditional security.
Let’s revise in some more detail the basic principle of both quantum cryptography and post-quantum cryptography.
4.1 Quantum Key Distribution: Classical cryptography is endangered by the discovery of the Shor’s algorithm because it can efficiently solve computationally hard problems upon which classical key-exchange mechanism are based. By contrast Quantum Key Distribution (QKD) does not make use of computationally hard primitives, but relies on the fundamental laws of quantum physics to establish security. However, it is worth to notice that QKD protocols are always hybrid, i.e., they rely on both quantum and classical communications to implement a virtually impenetrable crypto-system, promising to protect the privacy of communication even against attacks conducted by any quantum computers, independently from their computational power and evolution.
QKD can be implemented in two specific setups: continuous-variable (CV-QKD) and discrete-variable (DV-QKD) [43]. DV-QKD uses qubits to encode information and single photon detectors are employed by the receiver to monitor and quantify the presence of eavesdropper on the communication channel [45]. In such a way the parties can quantify the amount of information eavesdropped. By contrast, CV-QKD encodes classical information randomly modulating phase and amplitude of bright coherent states, and uses homodyne detection schemes at the receivers, in a similar setup used today by conventional optical communications [46].
@@
An essential tenet of QKD (both for DV and CV) is rooted in quantum physics and takes the shape of the no-cloning theorem [47, 48], which asserts that a flawless replica of arbitrary (i.e. non-orthogonal) quantum states cannot be created, without corrupting the probed quantum states. That is exploited during the quantum communication phase, when quantum particles are sent from the sender to the receiver. In fact, encoding information on non-orthogonal quantum states, makes so that any effort to gain insights on the properties of such a stream quantum signals would results in the introduction of noise, readily identified by either the key distributor or the recipient (the parties, conventionally Alice and Bob). Such a mechanism allows the parties to quantify the amount of information potentially eavesdropped (Eve) during the quantum communication. That information is crucial, because they can use it to then apply classical protocols of error correction and privacy amplification and reducing to a negligible amount the eavesdropper’s knowledge on the shared key. This second part of the cryptosystem is usually called classical communication phase. Examples of QKD protocols based on the steps described above are BB84, B92 and BM92 [49, 50, 51] that use DV-QKD, and CV-QKD protocols like [52, 53, 54].
Previous QKD protocols suffer from the relative vulnerabilities connected to imperfections and trust-ability of devices used in practical implementations. To overcome this difficulty, and potential security threats, an even more powerful approach to QKD has been introduced based on entanglement verification, and taking the name of Device-Independent (DI) QKD. In this approach the verification of violation of Bell inequalities is used to verify the presence of entanglement between the quantum signals shared between the parties. If entanglement is present then the parties will be in the position to share an unconditionally secure sequence of bits, ruling out any possibility for Eve to acquire information on the secret-key. The first seminal work using entanglement to implement QKD has been proposed by Ekert in his 1991 work [55]. After that many other works followed with refined security proofs [43].
DI-QKD is the ultimate approach to establish unconditionally secure secret-key using quantum mechanics without having to specify the physical implementation of equipments or fixing many potential quantum hacking loopholes [56]. However, DI-QKD is difficult to implement and its performance, on a practical scenario, are still limited [43] because it requires loophole-free Bell inequality violations, which necessitate high-quality entanglement among distant parties and near-perfect quantum detection, something current technologies cannot still provide in full [57], or at least not under commonly accepted practicality assumptions.
In recent years also Measurement Device-Independent (MDI) QKD protocols have been proposed to implement overcome difficulties connected to the trustability of measurement devices [58, 59, 60], and Twin-field QKD [61] to overcome the point-to-point quantum secret-key capacity, set by the PLOB bound [62], and recover the single-repeater scaling of end-to-end quantum capacity [63], without the need to implement a full-scale quantum repeater.
@@
4.2 Post-Quantum Cryptography: The security of classical cryptographic primitives (e.g., RSA, Diffie–Hellman, etc.) depends on the hard problems of discrete arithmetic, prime factorization of integers, and elliptic-curve discrete logarithms. Sadly, these present-day cryptographic primitives based on such hard problems might theoretically be solvable in a brief span of time using the possible applications of quantum computers. The potential attacks performed by quantum algorithms posed on conventional cryptographic protocols, have promoted a sense of urgency in designing alternative schemes to mitigate quantum attacks. Such alternatives are generally characterized as post-quantum cryptography (PQC). These schemes can effectively deal with prevalent challenges triggered by quantum adversaries. The threat represented by the potential implementation nof fast quantum algorithm able to break the conventional algorithm used in our everyday life lead to an intense research activity on identifying candidates algorithm for the implementation and update of communication infrastructure able to resits to attacks performed to know quantum algorithms  [64]. The protocols developed in post-quantum cryptography were generally grouped into five types: code-based, hash-based, lattice-based, multivariate, and supersingular curve-elliptic isogeny schemes [11].
NIST post-quantum cryptography standardization process [65] is underway to identify the specific algorithm families and protocols to be considered secure under the potential threat of a quantum computer.
It is worth to notice that the ultimate counter-measures to preserve security and privacy of communication against quantum eavesdroppers is quantum-key distribution (QKD), also against the possibility of “harvest now, decrypt later” approach. In which case attackers store encrypted material until when advances on decryption technology (hardware or software) allows to decrypt the stored content. It is clear that in case of extremely sensitive data this may represent a threat to security that cannot be neglected, i.e., where data needs to remain confidentially protected for very long period of time.